Consider the decomposition of a metal oxide to its
Answer #1

1)
delta Go = 2*delta Gof (M) + 1/2*delta Gof (O2) - delta Gof(M2O)
= 0 + 0 - (-7.90 KJ/mol)
= 7.90 KJ/mol

2)
delta Go = 7900 J/mol
use:
delta Go = -R*T*ln Kc
7900 = -8.314*298*ln Kc
Kc = 0.0412

3)
Kc= Kp*(RT)^delta n
0.0412 = kp*(0.0821*298)^(1/2)
kp = 8.33*10^-3

Kp = (p(O2))^1/2
p(O2) = kp^2
= (8.33*10^-3)^2
=6.94*10^-5 atm